OpenStudy (anonymous):

The function f(x) = 16(2)^x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (anonymous):

A) f(x) = 16(2)^x; growth rate 200% B) f(x) = 16(2)^2x; growth rate 8% C) f(x) = 16(1.41)^x; growth rate 8% D) f(x) = 16(1.41)^2x; growth rate 41%

OpenStudy (dan815):